# Matérn Kernels for Tunable Implicit Surface Reconstruction

### v1

@@ 4.1 Eigenvalue Decay, Spectral Bias, and Reconstruction Error
Eigenvalue decay and spectral bias. We begin by analyzing Matérn kernel’s eigenvalue decay rate (EDR) and show that it can be tuned by varying the smoothness parameter $\nu$ and/or bandwidth $h$. As investigated in Tancik et al. (2020) for coordinate-based MLPs, a rapid decrease of the associated NTK’s eigenvalues implies slow convergence to high-frequency components of the target function (up to the point where the network is practically unable to learn these components). Consequently, a smaller EDR, i.e., a slower eigenvalue decay, leads to faster convergence to high-frequency content; hence, in the context of implicit surface reconstruction, more detailed geometry. To overcome this slow convergence, a phenomenon known as spectral bias, Tancik et al. (2020) propose to use a (random) Fourier feature mapping of the form $\gamma(x)=[\cos(2\pi Bx),\sin(2\pi Bx)]$ applied to the inputs $x\in\mathbb{R}^{d}$ before passing them to the MLP, effectively transforming the MLP’s NTK, $k_{\text{NTK}}$, into a stationary kernel with tunable bandwidth parameter, $k_{\text{NTK}}(\gamma(x)^{\top}\gamma(y))=k_{\text{NTK}}(k(x-y))=:k^{\prime}_{\text{NTK}}(x-y)$. When $B\in\mathbb{R}^{q\times d}$ is drawn from a standard normal distribution (i.e., zero mean and unit variance), $k$ is an approximation of the Gaussian kernel (Rahimi & Recht, 2007). This allows explicit manipulation of the kernels’ spectrum, overcoming spectral bias.
@@
Figure 3 visualizes the EDR for different values of $\nu$ and $h$ along with the EDR of the non-tunable, first-order arc-cosine kernel which equates to $\Theta(s^{-(1+d)/d})$ and has only been very recently proven by Li et al. (2024) for general input domains and distributions. As seen, the smoother the kernel (i.e., the higher $\nu$) and the higher $h$, the faster the eigenvalues decay. This implies less detailed surface reconstruction (with a higher error) for smoother kernels and greater $h$; indeed, that is exactly what we observe in practice, see Figure 4. Contrary, by lowering both parameters, one can achieve faster convergence to high-frequency geometric details, effectively leading to more nuanced surface reconstruction and a smaller reconstruction error. Interestingly, the first-order arc-cosine kernel and Matérn 1/2 have the same EDR if $h=1$.
Figure 4: Matérn kernels are tunable. Surface reconstructions can be improved in practice by varying the bandwidth parameter, $h$, effectively tuning the kernels’ EDR (see also Figure 3). However, setting $h$ too small ($<0.5$) results in over-fitting, while setting $h$ too big ($>10$) over-smooths (i.e., under-fits) the true surface. This is also reflected in the reconstruction error (measured using Chamfer distance), plotted as a function of $h$ on the right.
Reconstruction error. We proceed by investigating the bandwidth’s influence on the reconstruction error, measured by the mean-squared error which is defined (and can be bounded) as:
@@
where $C_{\mathcal{X},\Omega}$ is a constant that only depends on $\mathcal{X}$ and $\Omega$ and for $\nu<\infty$. For more information on the bound, including details on $C_{\mathcal{X},\Omega}$, please see, e.g., Santin & Schaback (2016). By inspecting Eq. (12), we notice that the magnitude of the RKHS norm has a considerably high influence on the reconstruction error. Based on this observation and the fact that $\|f\|_{H_{\nu}}$ depends on $h$, our goal is now to further study the effect of $h$ on the RKHS norm. As a first result, we have:
@@
A proof can be found in Appendix A. We immediately see that $\|f\|^{2}_{H_{\nu}}\rightarrow\infty$ in either cases, $h\rightarrow 0$ and $h\rightarrow\infty$. Moreover, based on Proposition 2, it is easy to show that the norm, as a function of $h$, decreases as $h\rightarrow h^{*}$ from the left, and increases as $h\rightarrow\infty$, where $h^{*}$ is given as
@@
(see Appendix B for a derivation). This can also be seen from the inset figure on the right, where we plot the RKHS norm with respect to the obtained bound as a function of $h$.
Interestingly, our analysis reveals that the reconstruction error can not be made arbitrarily small by just lowering $h$; if the bandwidth $h$ is chosen too small, the reconstruction error starts to increase again. Conversely, if $h$ is set too high, the resulting surface is too smooth, ultimately leading to increased errors. In fact, what we actually observe is that Matérn kernels under-fit the true surface when $h$ is set too big, and, conversely, over-fit if $h$ is chosen too small, see also Figure 4. As we have shown, the optimal trade-off is reached if $h$ equals $h^{*}$. Moreover, whilst the bandwidth $h$ does play a critical role, our results show it is not very sensitive in practice, especially for Matérn 1/2 and 3/2 (see inset figure).
@@ 4.2 Relation to Neural Networks and the Arc-cosine Kernel
Relation to neural networks. Next, we study the connection between Matérn kernels and neural networks. It is well known that first-order arc-cosine kernels mimic the computation of a two-layer, infinite-width ReLU network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$, $f(x)=m^{-1/2}\sum_{i=1}^{m}v_{i}[w^{\top}_{i}x+b_{i}]_{+}$ when the bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$ are fixed from initialization and drawn from a standard normal distribution, see, e.g., Cho & Saul (2009); Williams et al. (2021). Here, $[x]_{+}:=\max\{0,x\}$ denotes the ReLU activation function.
Under similar assumptions, we now show that Matérn kernels mimic the computation of two-layer Sirens Sitzmann et al. (2020) (fully-connected networks with sine activation function) if the layer width approaches infinity. We claim the following:
@@
Consider a two-layer fully-connected network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ with sine activation function, $m$ hidden neurons, and fixed bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$. Let $h,\nu>0$ be given. If $w_{i}$ is randomly initialized from
@@
A proof can be found in Appendix C. This result establishes for the first time a connection between Matérn kernels and the computation in neural networks, in particular, Sirens.
Matérn kernels vs. arc-cosine kernels. Lastly, we aim to compare Matérn kernels with the previously used first-order arc-cosine kernels. Our analysis is based on:
@@
In other words, Theorem 4 shows that the RKHS of the Laplace kernel (a Matérn kernel with smoothness $\nu=1/2$) is the same as the RKHS induced by the NTK of a fully-connected ReLU network when inputs are restricted to $\mathbb{S}^{d-1}$. This immediately raises the question of whether the NTK for general input domains is significantly different from a standard Laplace kernel. Indeed, empirical results (Belkin et al., 2018; Hui et al., 2019; Geifman et al., 2020) already suggest a negative answer; however, from a theoretical perspective, the claim is still unproven. Based on Theorem 4, we obtain the following connection between Laplace and the first-order arc-cosine kernel:
@@
See Appendix D for a proof. For general input domains, our empirical results presented next support the claim mentioned above (at least for ($L=2$)-layer networks), showing that the Laplace kernel even outperforms the arc-cosine kernel.
@@ 5 Experiments and Results
We systematically evaluated the effectiveness of Matérn kernels in the context of implicit surface reconstruction, presenting results on ShapeNet (Chang et al., 2015) and the Surface Reconstruction Benchmark (SRB; Berger et al. (2013)) in Section 5.1. For ShapeNet, we use the train/val/test split provided by Mescheder et al. (2019). Leveraging Neural Kernel Fields (Williams et al., 2022), we present and evaluate learnable Matérn kernels in Section 5.2.
@@ 5.1 Surface Reconstruction on ShapeNet and SRB
ShapeNet. We compare Matérn kernels in a sparse setting against the arc-cosine kernel as implemented in Neural Splines (NS, Williams et al. (2021)). To do so, we randomly sample $m=1,000$ surface points with corresponding normals for each shape. We implemented Matérn kernels in PyTorch and took the official CUDA implementation of the arc-cosine kernel from NS, eventually integrated into a unified framework to ensure fair comparison. Notably, we did not use Nyström sampling as in NS, and employ PyTorch’s direct Cholesky solver to solve the KRR problem in Eq. (1) instead of an iterative conjugate gradient solver. Moreover, similar to NS, we approximate the gradient part in Eq. (1) with finite differences. See Appendix E for details.
F-Score $\uparrow$ CD $\downarrow$ 0.5 1 2 0.5 1 2 Matérn 1/2 93.6 93.7 93.7 4.05 4.02 4.02 Matérn 3/2 94.6 94.8 94.9 4.05 4.00 4.09 Matérn 5/2 92.4 93.8 92.9 6.42 5.65 6.91 Matérn $\infty$ 59.3 52.4 40.9 57.56 50.59 48.38 Arc-cosine 92.8 4.67 NS 90.6 4.74 SPSR 84.3 6.26       CD $\downarrow$ HD $\downarrow$ Time Matérn 1/2 0.21 4.43 9.85 sec Matérn 3/2 0.18 2.93 12.47 sec Matérn 5/2 0.58 22.52 15.26 sec Matérn $\infty$ 3.83 33.27 12.76 sec NS* 0.19 3.19 19.89 sec NS 0.17 2.85 11.91 sec SPSR 0.21 4.69 1.65 sec FFN 0.28 4.45 337.91 sec Siren 0.19 3.86 599.00 sec
@@
Quantitative results in terms of F-Score (with a threshold of 0.01) and Chamfer distance (CD; always reported $\times$ 103) can be found in Table 1. Matérn 1/2 and 3/2 consistently outperform the arc-cosine kernel as well as popular SPSR (Kazhdan & Hoppe, 2013). In addition, our experiments show that the bandwidth parameter, $h$, can be used to tune the surface reconstruction (lower the error), while not being too sensitive in practice. Matérn 5/2 and the Gaussian kernel (for $\nu\rightarrow\infty$) perform significantly worse, being simply too smooth. Figure 2 and Appendix E shows qualitative results.
SRB. Additionally, we evaluate Matérn kernels on the challenging Surface Reconstruction Benchmark (Berger et al., 2013) which consists of five complex shapes simulated from noisy range scans. For this, we integrated Matérn kernels into the NS framework, using KeOps (Charlier et al., 2021) and the FALKON (Rudi et al., 2017) solver. We used the same parameter settings as in the original NS paper. In addition, for Matérn kernels, we did a modest parameter sweep over the bandwidth, $h\in\{0.5,1,2\}$, and took the reconstruction with the lowest Chamfer distance.
@@
Figure 5: Qualitative results for non-learnable kernels on SRB (left) and learned kernels on ShapeNet (right). Please see Appendix F and G for more qualitative results.
@@
Quantitative results obtained by employing Chamfer and Hausdorff distance (HD) can be found in Table 1, demonstrating that Matérn 3/2 performs on par with NS while being faster to compute (compared to what we measured for NS). See Figure 5 (left) and Appendix F for qualitative results and per-object metrics.

### v2

@@ 3.2 Eigenvalue Decay, Spectral Bias, and Reconstruction Error
Eigenvalue decay and spectral bias. We begin by showing that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings help coordinate-based MLPs overcome spectral bias. As investigated in Tancik et al. (2020), a rapid decrease of the NTK’s eigenvalues implies an associated MLP’s slow convergence to high-frequency components of the target function (up to the point where the network is practically unable to learn these components). Consequently, a smaller eigenvalue decay rate (EDR), i.e., a slower eigenvalue decay, leads to faster convergence to high-frequency content—in the context of implicit surface reconstruction, more detailed geometry. To overcome this slow convergence, a phenomenon known as spectral bias, Tancik et al. (2020) use a Fourier feature mapping of the form
$\gamma(x)=[a_{1}\cos(2\pi b_{1}^{\top}x),a_{1}\sin(2\pi b_{1}^{\top}x),\dots,a_{q}\cos(2\pi b_{q}^{\top}x),a_{q}\sin(2\pi b_{q}^{\top}x)]^{\top}$ (9)
applied to the inputs $x\in\mathbb{R}^{d}$ before passing them to the MLP, effectively transforming the MLP’s NTK, $k_{\text{NTK}}$, into a stationary kernel, $k_{\text{NTK}}(\gamma(x)^{\top}\gamma(y))=k_{\text{NTK}}(k_{\gamma}(x-y))=:k^{\prime}_{\text{NTK}}(x-y)$ whose spectrum can be tuned through manipulation of the Fourier basis frequencies, $b_{i}\in\mathbb{R}^{d}$, and corresponding coefficients, $a_{i}\in\mathbb{R}$, of the kernel $k_{\gamma}(x-y)=\sum_{i=1}^{q}a^{2}_{i}\cos(2\pi b^{\top}_{i}(x-y))$. This shows that $\gamma$ enables explicit control over the NTK’s EDR, overcoming spectral bias.
@@
Figure 3 visualizes the EDR for different values of $\nu$ and $h$ along with the EDR of the non-tunable, first-order arc-cosine kernel which equates to $\Theta(s^{-(1+d)/d})$ and has only been very recently proven by Li et al. (2024) for general input domains and distributions. As seen, the smoother the kernel (i.e., the larger $\nu$) and the greater $h$, the faster the eigenvalues decay. This implies less detailed surface reconstruction (with a higher error) for smoother kernels and larger $h$; indeed, this is exactly what we observe in practice, see Figure 4. Contrary, by lowering both parameters, one can achieve faster convergence to high-frequency geometric details, effectively leading to more nuanced surface reconstruction and a smaller reconstruction error.
Figure 4: Matérn kernels are tunable. Surface reconstructions can be improved in practice by varying the bandwidth parameter, $h$, effectively tuning the kernels’ EDR (see also Figure 3). However, setting $h$ too small ($<0.5$) results in overfitting, while setting $h$ too big ($>10$) oversmooths (i.e., underfits) the true surface. This is also reflected in the reconstruction error (measured using Chamfer distance), plotted as a function of $h$ on the right.
Reconstruction error. We proceed by investigating the bandwidth’s influence on the $L_{2}$ reconstruction error, which is defined (and can be bounded) as
@@
where $C_{\mathcal{X},\Omega}$ is a constant that only depends on $\mathcal{X}$ and $\Omega$, and
$\|f\|_{H_{\nu}}^{2}=h^{2\nu}\left((2\pi)^{d/2}C_{d,\nu}\right)^{-1}\int_{\mathbb{R}^{d}}\left(\frac{2\nu}{h^{2}}+\left(2\pi\|\omega\|\right)^{2}\right)^{\nu+d/2}|\mathcal{F}[f](\omega)|^{2}\mathrm{d}\omega$ (11)
is the norm of the RKHS induced by a Matérn kernel with bandwidth $h$ and smoothness $\nu$. Moreover, $C_{d,\nu}:=(2^{d}\pi^{d/2}\Gamma(\nu+d/2)(2\nu)^{\nu})/\Gamma(\nu)$, and $\mathcal{F}[f]$ denotes the Fourier transform of a function $f$. For more information on the bound, including details on $C_{\mathcal{X},\Omega}$, please see, e.g., Santin & Schaback (2016). By inspecting Eq. (10), we notice that the magnitude of the RKHS norm has a considerably high influence on the reconstruction error. Based on this observation and the fact that $\|f\|_{H_{\nu}}$ depends on $h$, our goal is now to further study the effect of $h$ on the RKHS norm. As a first result, we have:
@@
A proof can be found in Appendix B. We immediately see that $\|f\|^{2}_{H_{\nu}}\rightarrow\infty$ in either cases, $h\rightarrow 0$ and $h\rightarrow\infty$. Moreover, based on Proposition 2, it is easy to show that the norm, as a function of $h$, decreases as $h\rightarrow h^{*}$ from the left, and increases as $h\rightarrow\infty$, where $h^{*}$ is given as
@@
Figure 5: RKHS norm as a function of $h$. We plot the bound from Proposition 2 as a function of $h$.
(see Appendix C for details). Our analysis shows that the reconstruction error can not be made arbitrarily small by just lowering $h$; if $h$ is chosen too small, the reconstruction error starts to increase again, i.e., it overfits the true surface. Conversely, if $h$ is set too high, the resulting surface is too smooth (underfits the true surface), ultimately leading to increased errors. The optimal trade-off is reached if $h=h^{*}$. This is also observed in practice, see Figure 4. As seen from Figure 5, however, the described effect is more pronounced for smoother Matérn kernels (large $\nu$), as, at some point, the norm (hence, the reconstruction error) increases rapidly for large $h$. Our theoretical analysis reveals several practical insights. First, the reconstruction error is generally more sensitive to very small values of $h$. Second, a good starting point is always $h=1$.
@@ 3.3 Relation to Neural Networks and the Arc-cosine Kernel
Relation to neural networks. Next, we study the connection between Matérn kernels and neural networks. It is well known that the first-order arc-cosine kernel mimics the computation of a two-layer, infinite-width ReLU network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$, $f(x)=m^{-1/2}\sum_{i=1}^{m}v_{i}[w^{\top}_{i}x+b_{i}]_{+}$ when the bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$ are fixed from initialization and drawn from a standard normal distribution, see, e.g., Cho & Saul (2009); Williams et al. (2021). Here, $[x]_{+}:=\max\{0,x\}$ denotes the ReLU activation function.
Under similar assumptions, we now show that Matérn kernels mimic the computation of two-layer Sirens (Sitzmann et al., 2020) if the layer-width approaches infinity. We claim the following:
@@
Consider a two-layer fully-connected network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ with sine activation function, $m$ hidden neurons, and fixed bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$. Let $h,\nu>0$ and $C_{d,\nu}$ be defined as in Eq. (11). If $w_{i}$ is randomly initialized from
@@
A proof can be found in Appendix D. This result establishes for the first time a connection between Sirens and kernel methods. While explicitly shown for Matérn kernels, as we argue in the Appendix D, the arguments of Theorem 3 in fact apply to all stationary kernels, making it a powerful tool to study the connection between widely used Sirens and any stationary kernel function.
Matérn kernels vs. arc-cosine kernels. Lastly, we aim to compare Matérn kernels with the previously used first-order arc-cosine kernel. Our analysis is based on:
@@
In other words, Theorem 4 shows that the RKHS of the Laplace kernel (a Matérn kernel with smoothness $\nu=1/2$) is the same as the RKHS induced by the NTK of a fully-connected ReLU network when inputs are restricted to $\mathbb{S}^{d-1}$. Based on Theorem 4, we obtain the following connection between Laplace and the first-order arc-cosine kernel:
@@
See Appendix E for a proof. For general input domains, our empirical results presented next support (at least for ($L=2$)-layer networks) the widely believed claim that the NTK is not significantly different from standard kernels (Belkin et al., 2018; Hui et al., 2019; Geifman et al., 2020), such as the Laplace kernel. Indeed, we show that the Laplace kernel even outperforms the arc-cosine kernel.
@@ 4 Experiments and Results
We systematically evaluated the effectiveness of Matérn kernels in the context of implicit surface reconstruction, presenting results on ShapeNet (Chang et al., 2015) and the Surface Reconstruction Benchmark (SRB; Berger et al. (2013)) in Section 4.1. Moreover, Section 4.2 demonstrates Matérn kernels’ ability to reconstruct high-frequency textures on the Google Scanned Objects (GSO; Downs et al. (2022)) dataset and Objaverse (Deitke et al., 2023). Leveraging Neural Kernel Fields (Williams et al., 2022), we present and evaluate learnable Matérn kernels in Section 4.3.
@@ 4.1 Surface Reconstruction on ShapeNet and SRB
ShapeNet. We compare Matérn kernels in a sparse setting against the arc-cosine kernel on ShapeNet (using train/val/test split provided by Mescheder et al. (2019)). To do so, we randomly sample $m=1,000$ surface points with corresponding normals for each shape. We implemented Matérn kernels in PyTorch and took the official CUDA implementation of the arc-cosine kernel from NS, eventually integrated into a unified framework to ensure fair comparison. Notably, we did not use Nyström sampling as in NS and employed PyTorch’s direct Cholesky solver to solve the KRR problem in Eq. (1) instead of an iterative conjugate gradient solver. Moreover, similar to NS, we approximate the gradient part in Eq. (1) with finite differences. See Appendix F for details.
F-Score $\uparrow$ CD $\downarrow$ 0.5 1 2 0.5 1 2 Matérn 1/2 93.6 93.7 93.7 4.05 4.02 4.02 Matérn 3/2 94.6 94.8 94.9 4.05 4.00 4.09 Matérn 5/2 92.4 93.8 92.9 6.42 5.65 6.91 Matérn $\infty$ 59.3 52.4 40.9 57.56 50.59 48.38 Arc-cosine 92.8 4.67 NS 90.6 4.74 SPSR 84.3 6.26      CD $\downarrow$ HD $\downarrow$ Matérn 1/2 0.21 4.43 Matérn 3/2 0.18 2.93 Matérn 5/2 0.58 22.52 Matérn $\infty$ 3.83 33.27 NS* 0.19 3.19 NS 0.17 2.85 SPSR 0.21 4.69 FFN 0.28 4.45 Siren 0.19 3.86 SAP 0.21 4.85 DiGS 0.18 3.55 VisCo 0.18 2.95 OG-INR 0.20 4.06
@@
Quantitative results in terms of F-Score (with a threshold of 0.01) and Chamfer distance (CD; always reported $\times$ 103) can be found in Table 1. Matérn 1/2 and 3/2 consistently outperform the arc-cosine kernel as well as popular SPSR (Kazhdan & Hoppe, 2013). In addition, our experiments show that the bandwidth parameter, $h$, can be used to tune the surface reconstruction (lower the error), while not being too sensitive in practice. Matérn 5/2 and the Gaussian kernel (for $\nu\rightarrow\infty$) perform significantly worse, being simply too smooth. Figure 2 and Appendix F shows qualitative results.
Time
Matérn 1/2 9.85
Matérn 3/2 12.47
Matérn 5/2 15.26
Matérn $\infty$ 12.76
NS* 19.89
NS 11.91
SPSR 1.65
Table 2: Runtime comparison on SRB. Time in seconds.
SRB. Next, we evaluate Matérn kernels on the challenging Surface Reconstruction Benchmark which consists of five complex shapes simulated from incomplete and noisy range scans with up to 100,000 points. For this, we implemented a highly scalable version of Matérn kernels, closely following the NS framework to solve the KRR problem in Eq. (1) using Nyström sampling (with 15,000 samples) and the FALKON (Rudi et al., 2017) solver. We compare against NS and SPSR, as well as kernel-free methods, Fourier Feature Networks (FFNs; Tancik et al. (2020)), Siren, SAP (Peng et al., 2021), DiGS (Ben-Shabat et al., 2022), VisCo (Pumarola et al., 2022), and OG-INR (Koneputugodage et al., 2023). All methods are optimization-based (FFN and Siren are overfitted to a single shape) and use surface normals. For Matérn kernels, we did a modest parameter sweep over $h\in\{0.5,1,2\}$ and took the reconstruction with the lowest Chamfer distance. Runtime is measured on a single NVIDIA V100.
@@
Figure 6: Qualitative results for non-learnable kernels on SRB (left) and learned kernels on ShapeNet (right). Please see Appendices G and I for more qualitative results.
@@
Quantitative results obtained by employing Chamfer and Hausdorff distance (HD) can be found in Table 1, demonstrating that Matérn 3/2 performs on par with NS while being significantly faster to compute (compared to what we measured for NS), see Table 2. Moreover, Matérn 3/2 outperforms all evaluated state-of-the-art kernel-free methods (FFN, Siren, SAP, DiGS, VisCo, and OG-INR). Figure 6 (left) and Appendix G provide more qualitative results and per-object metrics.
4.2 Texture Reconstruction on GSO and Objaverse
$m$ Kernel PSNR $\uparrow$ LPIPS $\downarrow$ 10,000 Matérn 3/2 19.61 2.05 Arc-cosine 19.34 2.07 2,500 Matérn 3/2 18.92 2.15 Arc-cosine 18.88 2.09
Table 3: Texture reconstruction on GSO. Matérn kernels outperform the arc-cosine kernel in dense and sparse settings.
Lastly, we demonstrate Matérn kernels’ ability to represent other high-frequency scene attributes, such as texture. To do so, we randomly sample $m$ surface points with corresponding normals and per-point RGB color values from textured meshes, yielding an extended dataset $\mathcal{D}^{\prime}=\mathcal{D}\times\{c_{1},c_{2},\dots,c_{m}\}\subset\Omega\times\mathbb{R}^{d}\times\mathbb{R}^{3}$ for each shape. Then, instead of modeling an SDF $f$ as in Eq. (1), we are seeking a function $f^{\prime}:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{4}$ that, along with signed distances, also predicts per-point RGB values. Please find more information in Appendix H about how we adapt the KRR problem in Eq. (1) to estimate $\hat{f}^{\prime}$. Finally, we extract the object’s surface using Marching Cubes and trilinearily interpolate RGB values at previously predicted surface points.
Figure 7: Texture reconstruction on Objaverse. Matérn kernels lead to fewer artifacts and sharper details than the arc-cosine kernel.
We present quantitative results in terms of PSNR and LPIPS on GSO in Table 3 (metrics evaluated on the texture atlas; see Appendix H for further details). Matérn 3/2 surpasses the arc-cosine kernel in the densely sampled setting using 10,000 surface points, as well as in the sparse setting. Notably, we did not tune Matérn kernels’ bandwidth parameter for this experiment.
We also show qualitative results on the challenging Objaverse (Deitke et al., 2023) dataset in Figure 7. We chose Objaverse as it includes extremely high-resolution and complex textures. Matérn 3/2 reconstructs high-frequency texture details with great precision, overall yielding visually more pleasant reconstructions than the arc-cosine kernel. Reconstructed textures have fewer artifacts and are generally sharper, demonstrating that Matérn kernels can overcome spectral bias much better than the arc-cosine kernel. Note that this perfectly confirms our theoretical analysis presented in Section 3.2.